### v1

@@ Discussion
A key bottleneck in this study is the exposure time of our RGB camera (2s per snapshot), as SPARX’s inference itself takes only fractions of a second (Figure 5(b)). In our experiment, long exposure times–2s here in Figure 5(a)– without gain are used to achieve high-quality Airy patterns with minimal noise, since no significant spectral peaks are captured by the RGB channels below 700 nm, as shown in SI Fig. S3(a). Multiple strategies can be implemented to improve throughput without significantly compromising spectral prediction accuracy. Most naively, expanding the FOV with lower-magnification objectives or wide-field illumination designs could capture larger sample areas per snapshot, increasing the number of nanoparticles analyzable per cycle. From the DL model’s point of view, while increasing the CMOS camera gain introduces stochastic noise, DL models usually demonstrate inherent noise robustness, maintaining prediction accuracy despite elevated noise levels. This enables the use of shorter exposure times, thus effectively improving the acquisition speed. Further improvement could be achieved by augmenting the illumination intensity. High-brightness light sources, such as supercontinuum white-light lasers or laser-driven light sources, could enhance scattering signals from nanoparticles, reducing the exposure time by several orders of magnitude required for sufficient DF image contrast.
Enhancing spatial resolution represents another critical pathway. Employing higher NA objectives coupled with highly confocal optical configurations would improve imaging resolution, enabling the precise characterization of densely packed nanoparticles. Integrating advanced imaging technologies, such as structured illumination microscopy [40], super-resolution plasmonics [41], or DF nanoscopy [42], could further bypass the diffraction limit, achieving nanoscale spatial resolution and facilitating high-density nanoparticle analysis.
@@
In addition, spectral optimization presents two complementary avenues. Firstly, one can tailor the plasmonic systems to exhibit dominant resonant peaks within the visible range (400–700 nm). It would align spectral signatures with the RGB camera’s detection window, maximizing signal capture efficiency and allowing significantly shorter exposure times. Secondly, extending detection beyond 700 nm using hyperspectral imaging [43] , e.g., multi-channel split-frequency detection systems (e.g., grayscale CCDs with spectral filters) could broaden the spectral window to 1000 nm or longer wavelengths, capturing near-infrared resonances.
Another key opportunity for improvement lies in refining the assumptions behind the heteroscedastic loss function. Currently, a Gaussian distribution is assumed for prediction errors across wavelengths—an approach that simplifies training and serves as a useful approximation. However, this assumption may not capture the true error structure in cases involving highly nonlinear or complex nanoparticle geometries. Exploring alternative statistical models, such as heavy-tailed or skewed distributions, could provide more accurate uncertainty estimation and enhance the robustness of predictions.
In parallel, optimizing the network architecture and exploring more advanced DL models could yield significant performance gains. Fine-tuning hyperparameters like layer depth, filter size, and learning rate, possibly through Bayesian optimization or genetic algorithms, may uncover more efficient configurations. Additionally, probabilistic models such as variational autoencoders (VAEs) combined with Monte Carlo sampling can capture the stochastic nature of spectral responses at the single-particle level. Bayesian neural networks (BNNs) also offer a principled way to model uncertainty by treating weights as distributions, potentially improving generalization and performance in data-limited scenarios.
@@ Conclusion
We developed the SPARX model, a deep learning-based paradigm for fast and accurate spectral extrapolation from information-limited RGB DF images, beyond the RGB camera’s capture capabilities. SPARX batch-predicts nanoparticle spectra with millisecond latency, achieving a throughput 3-4 orders of magnitude faster than traditional serial spectroscopic measurements, while maintaining accuracy comparable to direct acquisitions. Beyond prediction, SPARX quantifies uncertainty using a heteroscedastic model, providing an additional key for the reliable screening of high-confidence predictions. By combining accurate extrapolation, batch processing, and reliable uncertainty quantification, SPARX paves the way for high-throughput, reproducible optical characterizations. This opens new possibilities for robust single-particle level nanophotonic devices, biochemical sensing and imaging, and other multidisciplinary applications.
@@ Methods
The optical setup used to obtain the data for this study is illustrated in Figure 2(a). The DF imaging of NPoM systems was implemented using a commercial illuminator (Olympus BX53). A halogen lamp served as the white light source, collimated through a condenser lens and subsequently directed to a DF module containing a 45°-tilted annular mirror. This optical configuration generated annular illumination that was coupled into the outer annular channel of a DF objective (NA=0.9, working distance=1 mm), producing grazing-incidence excitation at the sample plane. Scattered light from individual NPoM nanostructures was collected through the same objective and subsequently split by a 50/50 beam splitter (Chroma). The reflected optical path was directed to an imaging module where a tube lens focused the signal onto an RGB CMOS camera (Tucsen MIchrome 20), enabling wide-field DF imaging. The transmitted path was coupled into a spectrometer (Horiba IHR320) through a motorized entrance slit, with spectral dispersion achieved via a grating (150 l/mm) and detection using a CCD. Spatially resolved single-particle spectroscopy was achieved through confocal alignment optimization, where the spectrometer slit width was precisely adjusted to match the dimension of individual nanoparticles. DF spectra were normalized using the expression $S=(A-B)/L$, where A represents the raw scattering spectrum, $B$ denotes the background spectrum acquired from adjacent mirror regions, and $L$ corresponds to the illumination source spectrum. An home-built automated measurement protocol was implemented for high-throughput characterization. Individual NPoM structures were localized via particle tracking algorithms and precisely positioned at the optical axis center using a motorized XYZ translation stage. Synchronized acquisition of both spectral and imaging data streams was achieved through custom control software, enabling correlated structural and optical analysis of single NPoM nanoantennas.
@@
$\mathcal{L}_{\text{NLL}}=\frac{1}{N}\sum_{i=1}^{N}\left[\log\sigma^{2}(\lambda_{i})+\frac{(y(\lambda_{i})-\mu(\lambda_{i}))^{2}}{\sigma^{2}(\lambda_{i})}\right]+\text{const.}$. (2)

### v2

@@ Discussion
A key bottleneck in Figure 5 is the exposure time of our RGB camera (2s per snapshot for high resolution), as SPARX’s inference itself takes only fractions of a second (Figure 5(b)). Future improvements in throughput and performance can be approached from both algorithmic and optical perspectives. Algorithmically, SPARX inference is already fast, but accuracy and uncertainty estimation could be enhanced by refining the heteroscedastic loss function, exploring alternative error distributions (e.g., heavy-tailed or skewed), optimizing network architectures and hyperparameters, or adopting advanced models such as variational autoencoders (VAEs) or Bayesian neural networks (BNNs) to better capture stochastic single-particle spectral responses. Optically, acquisition speed could be improved by increasing camera gain (leveraging the model’s noise robustness), expanding the field of view with lower-magnification objectives or wide-field illumination, boosting illumination intensity using high-brightness sources, and enhancing spatial resolution with higher NA objectives, confocal configurations, or advanced techniques such as structured illumination microscopy [60], super-resolution plasmonics [61], or DF nanoscopy [62]. These combined algorithmic and optical strategies can significantly accelerate measurements while maintaining or improving spectral prediction accuracy.
@@
From the perspective of information acquisition, spectral optimization presents two complementary avenues. Firstly, one can tailor the plasmonic systems to exhibit dominant resonant peaks within the visible range (400–700 nm). It would align spectral signatures with the RGB camera’s detection window, maximizing signal capture efficiency and allowing significantly shorter exposure times. Secondly, inspired by hyperspectral techniques [63], a few detection channels beyond 700 nm can be incorporated—for example, via multi-channel split-frequency systems such as grayscale CCDs with spectral filters—to broaden the spectral window up to 1000 nm or beyond, thereby capturing near-infrared resonances.
Beyond Spectral Prediction: Shape Classification from DF RGB Images
Beyond spectral reconstruction, DF RGB images encode additional information that SPARX can exploit for nanoparticle shape classification. Since metallic nanoparticle’s shape strongly influences plasmonic performance [64], and synthesis often yields by-products with varied geometries, this approach provides a valuable, noninvasive alternative to conventional techniques such as scanning electron microscopy (SEM), transmission electron microscopy (TEM), and atomic force microscopy (AFM). In future applications, SPARX classifier could distinguish multiple particle geometries simultaneously on the same substrate, enabling additional rapid screening of desired nanoparticles or studies that leverage spectral diversity from different particle types [65].
The underlying physical motivation for classification lies in the pronounced differences in the spatial distributions of near- and far-field among nanoparticles of different shapes (e.g., nanocubes and nanospheres shown here). These differences are expected to be reflected in their DF RGB images (SI Fig. S8(a)). Again, these RGB images can be challenging for human vision to distinguish, as visually similar images may belong to different classes. Before developing a supervised classification model, we first conducted an unsupervised study to explore the natural separability of the two datasets (SI Fig. S8. We normalized each RGB image to its maximum pixel value, removing the influence of acquisition time and variations in optical collection efficiency. This ensures that any clustering arises from intrinsic image features rather than overall brightness. We also plotted the intensity distributions for each RGB channel in Fig. S8(b), which display similar patterns. The UMAP distribution RGB images in Fig. S8(c) shows that the nanospheres are relatively separable from nanocubes. These findings support the feasibility of using a supervised model to achieve robust nanoparticle classification.
We first explored a PCA–LDA (Linear Discriminant Analysis) pipeline as a base model for distinguishing nanocubes from nanospheres solely using DF RGB images. As shown in Figs. 6(a) and (b), the classification accuracy improved with increasing PCA dimensionality, with a substantial rise between 10 and 30 components. Based on this, we evaluated three representative cases: 2, 10, and 30 PCA components, achieving accuracies of 75%, 91%, and 95%, respectively.
Figure 6: PCA–LDA classification of nanospheres and nanocubes from DF RGB images. (a) Classification accuracy as a function of the number of PCA components used before LDA, showing significant performance gains at 10 and 30 components. (b) The comparison of the accuracy of different models. (c) Reconstructed LDA projection vectors for 2, 10, and 30 PCA components, reshaped into image space to highlight spatial–spectral features most relevant for discrimination. (d) Distributions of inner-product scores between the projection vectors and normalized DF images for both particle types, with optimal decision thresholds indicated. Accuracies of 75%, 91%, and 95% were achieved for 2, 10, and 30 PCA components, respectively.
To gain physical insight into the classification process, we reconstructed the LDA discriminant vectors into image space. These projection vectors, shown in Fig. 6(c), highlight spatial and spectral regions most relevant for discrimination. For example, in the 30D PCA case of Fig. 6(c), the green channel is stronger at the image center, while the red channel is weaker. Combined with the kernel density estimation (KDE) plots in Fig. 6(d)—where spheres have higher projection scores than cubes—this indicates that, statistically, nanospheres exhibit higher green intensity at the center of their DF images, whereas nanocubes display more pronounced red in the same region. Physically, when the overall dimensions are similar, this difference may arise from the distinct far-field distributions of the two geometries: nanospheres tend to support more symmetric scattering modes that enhance shorter-wavelength (green) scattering at the center, while nanocubes, with sharper edges and corners, can localize fields in a way that favors longer-wavelength (red) scattering in the central region [59]. These projection vectors can also be used directly for classification by correlating them with normalized DF images and applying a single decision threshold (Fig. 6(d)), effectively combining feature extraction and classification in one interpretable step.
For comparison, we trained a SPARX classifier, whose architecture is shown in Fig. S9. As illustrated in Fig. 6 (a) with a red bar, this model achieved a classification accuracy of 99.8%, substantially higher than the best PCA–LDA results. One likely reason for this improvement is that SPARX classifier, as a convolutional neural network, is inherently more robust to variations in the spatial position of features within an image. In our case, this means that a DF image of a nanoparticle can be accurately classified even if its scattering pattern appears at a slightly different location in the frame. In contrast, methods such as LDA are sensitive to absolute pixel positions, and the same DF pattern shifted spatially within the image may be classified incorrectly.
In summary, we have demonstrated the SPARX framework, which reveals the rich information hidden in simple DF RGB images. From these RGB signals, SPARX can reconstruct the broadband spectra and, importantly, extract and understand information for specialized applications such as nanoparticle screening and shape classification. It is reasonable to anticipate that other properties related to DF scattering, such as particle size or nanogap thickness, could also be inferred directly from DF RGB images in the future. In contrast, conventional spectral acquisition methods, including HSI, can precisely capture spectra, yet they do not decode the underlying multimodal resonance correlations inherent to optical systems, limiting their suitability for inference tasks such as classification.
However, our DL-based SPARX reduces both cost and technical barriers: after training, it relies solely on standard dark-field microscopy components (e.g., an RGB camera) and does not require a spectrometer, whereas conventional hyperspectral systems typically demand more expensive, specialized hardware and stricter spectral–spatial calibration. This positions SPARX as an accessible, lab-ready solution for general optical characterization and specialized plasmonic applications—validated across diverse nanoplasmonic geometries, including nanospheres and nanocubes as demonstrated in this work. And SPARX is compatible with emerging techniques such as compressed sensing, enabling further acceleration when dealing with sparse samples.
@@ Conclusion
We developed the SPARX framework, a deep learning-based paradigm to rapidly and accurately extrapolate spectra from information-limited RGB DF images, classify the shape of the nanoparticles, beyond the RGB camera’s capture capabilities. SPARX batch-predicts nanoparticle spectra with millisecond latency, achieving a throughput 2-4 orders of magnitude faster than traditional serial spectroscopic measurements, while maintaining accuracy comparable to direct acquisitions. Beyond spectral prediction, SPARX quantifies uncertainty using a heteroscedastic model, providing an additional key for the reliable screening of high-confidence predictions. Furthermore, the SPARX classifier enables inference tasks for additional screening of the samples, which are challenging for conventional spectroscopic methods.
Therefore, SPARX opens exciting possibilities for addressing a critical challenge in extreme nanophotonics: the intrinsic trade-off between exceptional single-particle performance and lack of reproducibility across a population, a long-standing bottleneck. By combining accurate extrapolation, batch processing, reliable uncertainty quantification, inference capability, and transplantability across diverse nanogap geometries, SPARX paves the way for multi-task, high-throughput, and reproducible measurements. Beyond plasmonic applications, the general SPARX paradigm—replacing spectrometer-dependent workflows with camera-based deep learning—establishes a scalable and lab-ready methodology for next-generation optical characterization and technology.
@@ Methods
The optical setup used to obtain the data for this study is illustrated in Figure 2(a). The DF imaging of NPoM systems was implemented using a commercial illuminator (Olympus BX53). A halogen lamp served as the white light source, collimated through a condenser lens and subsequently directed to a DF module containing a 45°-tilted annular mirror. This optical configuration generated annular illumination that was coupled into the outer annular channel of a DF objective (NA=0.9, working distance=1 mm), producing grazing-incidence excitation at the sample plane. Scattered light from individual NPoM nanostructures was collected through the same objective and subsequently split by a 50/50 beam splitter (Chroma). The reflected optical path was directed to an imaging module where a tube lens focused the signal onto an RGB CMOS camera (Tucsen MIchrome 20), enabling wide-field DF imaging. The transmitted path was coupled into a spectrometer (Horiba IHR320, CCD: Symphony II FIVS) through a motorized entrance slit, with spectral dispersion achieved via a grating (150 l/mm) and detection using a CCD. Spatially resolved single-particle spectroscopy was achieved through confocal alignment optimization, where the spectrometer slit width was precisely adjusted to match the dimension of individual nanoparticles. DF spectra were normalized using the expression S=(AB)/LS=(A-B)/L, where A represents the raw scattering spectrum, $B$ denotes the background spectrum acquired from adjacent mirror regions, and $L$ corresponds to the illumination source spectrum. An home-built automated measurement protocol was implemented for high-throughput characterization. Individual NPoM structures were localized via particle tracking algorithms and precisely positioned at the optical axis center using a motorized XYZ translation stage. Synchronized acquisition of both spectral and imaging data streams was achieved through custom control software, enabling correlated structural and optical analysis of single NPoM nanoantennas.
@@
$\mathcal{L}_{\text{NLL}}=\frac{1}{N}\sum_{i=1}^{N}\left[\log\sigma^{2}(\lambda_{i})+\frac{(y(\lambda_{i})-\mu(\lambda_{i}))^{2}}{\sigma^{2}(\lambda_{i})}\right]+\text{const.}$ (2)
@@
PCA–LDA Classification Pipeline
For PCA–LDA classification, DF RGB images (originally 128×128 pixels) were resized to 32×32 pixels to reduce computational cost. Each image was then flattened from (32, 32, 3) to a vector of length 3072 before PCA transformation. PCA reduced the data to $n$ components, which were then passed to an LDA classifier. The number of PCA components was varied to study the effect on classification accuracy.
For discriminant visualization, the LDA projection vector was mapped back to the original data space by multiplying it with the PCA transformation matrix and then reshaping it to (32, 32, 3). These projection vectors were used both for visual interpretation and as direct classifiers. For direct classification, the inner product between the projection vector and a normalized DF image was computed, and a threshold was applied to determine the particle type. Score distributions for nanospheres and nanocubes, along with optimal thresholds, are provided in Fig. 6.